OpenStudy (anonymous):

What are the solutions to the quadratic equation 5x2 + 60x = 0? x = 0 and x = −12 x = 0 and x = 12 x = 5 and x = −12 x = 5 and x = 12

OpenStudy (taramgrant0543664):

Do you have an idea on how to do this?

OpenStudy (anonymous):

I think I do somewhat

OpenStudy (usukidoll):

factor out the like terms for \[5x^2+60x=0\]

OpenStudy (anonymous):

Aren't I supposed to find what 5^2 and 60 have in common or something like that or I am mistaken ._.

OpenStudy (usukidoll):

yeah.... you are supposed to find what they have in common and yank em out.

OpenStudy (taramgrant0543664):

Well in order to get 0 on the right we know that x=0 in one case because substituting 0 in we will get 0 on the left

OpenStudy (taramgrant0543664):

To find the second number you have to factor out what you can so what can you factor out?

OpenStudy (usukidoll):

let's split this up for a sec \[5x^2+(5)(12)x=0\] ok what number do they have in common and what variable do they have in common?

OpenStudy (anonymous):

They have 5 and 12 in common.... 5 and 12 have..... 3 in common I guess.... since 3 times 3 is 12 and 3 +2 = 5

OpenStudy (taramgrant0543664):

They both have 5x in common so you can factor that out

OpenStudy (taramgrant0543664):

So 5x(x+12)=0

OpenStudy (usukidoll):

how is 3 x 3 12? 3 x 3 = 9

OpenStudy (anonymous):

I meant 3 times 4 sorry

OpenStudy (taramgrant0543664):

To get the inside the brackets equal to 0 how do you get that?

OpenStudy (usukidoll):

well 12 x 5 = 60. when I split it up we have a 5 for the number in common and x for the variable in common so 5x gets yanked out and we're left with 5x(x+12)=0 now we solve for x for these two cases 5x = 0 and x +12 = 0

OpenStudy (anonymous):

By subtracting because we want less?

pooja195 (pooja195):

Solve both for x

OpenStudy (anonymous):

x = 0 its so simple ._.

OpenStudy (usukidoll):

@pooja195 yes solve both equations for x

pooja195 (pooja195):

5x = 0 (divide) x +12 = 0 (sibtract)

OpenStudy (taramgrant0543664):

(X+12)=0 you want some number to be negative so it equals 0

OpenStudy (usukidoll):

we want x by itself for x+12=0 so we subtract both sides by 12

OpenStudy (anonymous):

The answer is still zero since 0 minus 12 is 12 and 12 divided by 0 is ZERO so both answers are Zero :D

OpenStudy (usukidoll):

omg I am facepalming over this.

OpenStudy (usukidoll):

there are two solutions of x.

OpenStudy (usukidoll):

we got x = 0 for 5x=0 now solve x+12 = 0 by subtracting 12 on both sides.

OpenStudy (anonymous):

Its 11pm I have had a long day I am already bad at math add me doing this half asleep.....

OpenStudy (anonymous):

I DID!

pooja195 (pooja195):

0-12?

OpenStudy (taramgrant0543664):

X+12=0 x+12-12=0-12

OpenStudy (anonymous):

GOD ITS LIKE 6th grade ALGEBRA! 12 minus zero is 12

OpenStudy (taramgrant0543664):

Zero minus twelve not twelve minus zero

OpenStudy (anonymous):

Well Negative 12

OpenStudy (usukidoll):

thank you @taramgrant0543664

OpenStudy (anonymous):

But it is still 12

OpenStudy (anonymous):

Just on the left side XD

OpenStudy (taramgrant0543664):

Zero minus twelve is negative twelve

OpenStudy (anonymous):

I just said that ._.

OpenStudy (usukidoll):

5x(x+12) = 0 5x = 0 divide 5 on both sides and we have x = 0 x+12 = 0 -12 -12 x = -12

OpenStudy (taramgrant0543664):

It's not twelve it's negative twelve

OpenStudy (anonymous):

Well sorry! The answer is A i finally finished the test thanks and sorry for being a freaking idiot..... its just late and I have been doing work all day and by now my head is just on fire ._. I am going to bed good night

OpenStudy (anonymous):

@UsukiDoll Sarcasm is one of my many flaws its an aspect of my personality I have to fix...... I am very sorry if I was a jerk or I offended you in any way thanks for the help and patients both you and @taramgrant0543664

OpenStudy (usukidoll):

ah... that's ok. I take meds in real life or get high blood pressure so I type some crazy stuff sometimes too. XD

OpenStudy (mathstudent55):

\(\Large 5x^2 + 60x=0\) Factor out the common factor 5x: \(\Large \color{red}{5x} \cdot x + 12 \cdot \color{red}{5x} = 0\) \(\Large 5x(x + 12) = 0\) Divide both sides by 5: \(\Large x(x + 12) = 0\) Set both factors equal to 0: \(\Large x = 0\) or \(\Large x + 12 = 0\) Solve both equations for x: \(\Large x = 0\) or \(\Large x = -12\)

OpenStudy (anonymous):

lol ok

OpenStudy (usukidoll):

meow ^.^

OpenStudy (anonymous):

Oh and I got an 89% in the test! XD

OpenStudy (anonymous):

So double thanks now good night and bye